# Context-Aware Video Instance Segmentation

### v1

@@ 4.2 Prototypical Cross-frame Contrastive Loss
Our method employs a query-based segmenter to assess each pixel’s relationship to an object by comparing core instance features, $\hat{Q}_{t}^{n}\in\mathbb{R}^{C}$, with corresponding pixel embeddings, $F_{t}^{\{h,w\}}\in\mathbb{R}^{C}$. This process ensures consistent feature representation within object-containing regions and introduces an intra-frame constraint that reflects similarity within the feature map. By examining the inter-frame relationships of pixel embeddings, our method improves object association, essential for contrastive learning methods that strive to differentiate between similar and dissimilar objects across frames.
@@
$\mathcal{L}_{\text{PCC}}=\sum_{t=1}^{T}\sum_{n=1}^{N_{GT}}\mathcal{L}_{\text{Emb}}\left(\eta_{t}^{\hat{\sigma}^{t}(n)}\right),~{}~{}\eta_{t}^{n}=\frac{\sum_{h=1}^{H}\sum_{w=1}^{W}F_{t}^{\{h,w\}}*\mathbbm{1}\left(M_{t}^{\{h,w\}}==1\right)}{\sum_{h=1}^{H}\sum_{w=1}^{W}\mathbbm{1}\left(M_{t}^{\{h,w\}}==1\right)}.$ (11)
@@ 4.3 Training Loss
To train the tracking network $\mathcal{T}$, we calculate the matching cost only for objects that appear for the first time to ensure consistent video-level matching pairs, as implemented in prior work (Zhang et al., 2023a). The objective function $\mathcal{L}_{\mathcal{T}}$ incorporating the matching relation $\hat{\sigma}_{\text{con}}$ is defined as follows:
$\begin{gathered}\mathcal{L}_{\mathcal{T}}=\sum_{t=1}^{T}\sum_{n=1}^{N_{GT}}\mathcal{L}_{\text{Inst}}\left(y^{t}_{n},\hat{y}^{t}_{\hat{\sigma}_{\text{con}}(n)}\right),~{}~{}\hat{\sigma}_{\text{con}}=\underset{\sigma\in\mathfrak{S}_{N}}{\arg\min}\sum_{k=1}^{N_{GT}}\mathcal{L}_{\text{Match }}\left(y_{k}^{f(k)},\hat{y}_{\sigma(k)}^{f(k)}\right),\end{gathered}$ (13)
@@ 5 Experiments
We evaluate CAVIS on two major tasks: video instance segmentation (VIS) and video panoptic segmentation (VPS) on four benchmark datasets recognized for their challenges and prevalence in the research community: YouTubeVIS-2019 (Yang et al., 2019), YouTubeVIS-21 (Yang et al., 2021a), OVIS (Qi et al., 2022), and VIPSeg (Miao et al., 2021). For VIS, performance metrics include average precision (AP) and average recall (AR) as established in previous studies (Yang et al., 2019). In the realm of VPS (Kim et al., 2020), we further examine our model’s capabilities using the Segmentation and Tracking Quality (STQ) metric, and the Video Panoptic Quality (VPQ) metric.
Table 1: Comparisons on the validation sets of YouTube-VIS 2019, 2021, and OVIS datasets. The best and second-best scores are highlighted in red and blue, respectively. $\dagger$ denotes the model trained with the temporal refiner (Zhang et al., 2023a). Rows in cyan indicate comparisons with a top-performing model.
[…]
@@ 5.1 Implementation Details
We employ Mask2Former (Cheng et al., 2022) as our segmentation network, utilizing three distinct backbone encoders: ResNet-50 (He et al., 2016), Swin-L (Liu et al., 2021), and ViT-L (Dosovitskiy et al., 2021). The ResNet-50 and Swin-L backbones are initialized with parameters pre-trained on the COCO dataset (Lin et al., 2014), while the ViT-L backbone uses initialization parameters from DINOv2 (Oquab et al., 2023). Additionally, for the ViT-L backbone, we employ a memory-efficient version of VIT-Adapter (Chen et al., 2022), aligning with recent advancements in network efficiency (Zhang et al., 2023b). Further details are described in Sec. A.2.2.
@@ 5.2 Comparison to State-of-the-Art Methods
Figure 3: Qualitative comparisons of CAVIS (ours) against state-of-the-art methods: CTVIS (Ying et al., 2023) and DVIS++ (Zhang et al., 2023b) on the OVIS dataset.
@@ 5.3 Ablation Study
We conduct ablation studies on the OVIS dataset (Qi et al., 2022) with the ResNet-50 (He et al., 2016) backbone over 40k training iterations, detailed in Tab. 3(f). For these tests, we sample two frames from each input video for experiments listed in Tab. 3(f)-(a), (b), and (d). To evaluate the segmentation network across these setups in Tab. 3(f)-(a-d), we employ the minimal post-processing method proposed by MinVIS (Huang et al., 2022).
@@
Ablation study on technical contributions of CAVIS. We conduct a series of experiments to demonstrate the effectiveness of our key components: the context-aware instance tracker (CAIT) and prototypical cross-frame contrastive (PCC) loss. The results, shown in Tab. 3(f)-(a) for the segmentation network and Tab. 3(f)-(e) for the tracking network, demonstrate significant performance enhancements.
@@
Tab. 3(f)-(a) includes six experiments (i- vi), where experiment (i) present our baseline performance of retraining MinVIS (Huang et al., 2022) to match our settings. Experiments (i- iii) show that implementing contrastive learning, whether with standard or context-aware instance features, leads to significant performance gains. Particularly, context-aware instance features result in a notable +2.7 AP improvement over the baseline, a considerable increase compared to the +1.3 AP improvement observed with standard instance features. Further performance boosts are noted when PCC is introduced in experiments (iv- vi). The most substantial improvement is seen in experiment (vi), which records the highest performance of 29.5 AP by utilizing both $\mathcal{L}_{\text{CTX}}$ and $\mathcal{L}_{\text{PCC}}$. This highlights the synergistic effect of integrating context-aware tracking with cross-frame contrastive loss, significantly enhancing the system’s accuracy and effectiveness.
Table 2: Comparison on VIPSeg validation sets. ‘Th’ and ‘St’ denote ‘things’ and ‘stuff’ classes.
Method ResNet-50 ViT-L VPQ $\text{VPQ}^{\text{Th}}$ $\text{VPQ}^{\text{St}}$ STQ VPQ $\text{VPQ}^{\text{Th}}$ $\text{VPQ}^{\text{St}}$ STQ VPSNet-SiamTrack(Woo et al., 2021) 17.2 17.3 17.3 21.1 - - - - VIP-Deeplab(Qiao et al., 2021) 16.0 12.3 18.2 22.0 - - - - Clip-PanoFCN(Miao et al., 2022) 22.9 25.0 20.8 31.5 - - - - Video K-Net (Li et al., 2022) 26.1 - - 31.5 - - - - TarVIS(Athar et al., 2023) 33.5 39.2 28.5 43.1 - - - - Tube-Link(Li et al., 2023c) 39.2 - - 39.5 - - - - Video-kMax (Shin et al., 2024) 38.2 - - 39.9 - - - - DVIS (Zhang et al., 2023a) 39.4 38.6 40.1 36.3 - - - - DVIS++ (Zhang et al., 2023b) 41.9 41.0 42.7 38.5 56.0 58.0 54.3 49.8 Ours 42.4 43.1 41.8 39.7 56.9 60.1 54.2 51.0 DVIS $\dagger$ (Zhang et al., 2023a) 43.2 43.6 42.8 42.8 - - - - DVIS++ $\dagger$ (Zhang et al., 2023b) 44.2 44.5 43.9 43.6 58.0 61.2 55.2 56.0 Ours $\dagger$ 45.3 47.5 43.4 45.3 58.5 63.1 54.5 56.1
@@
In Tab. 3(f)-(e), we evaluate the tracking network by initially using a fixed pre-trained segmentation network. Employing standard cross-attention with $\hat{Q}$ achieved a high performance of 34.4 AP. Substituting this with our newly designed context-aware cross-attention further increases the performance by +1.7 AP, reaching 36.1 AP. These findings validate the efficacy of the context-aware feature, confirming its significant advantages for instance matching in complex video scenarios.
@@
Metric Context alignmnet
@@
Context filter. Given that images feature objects at various scales, identifying the optimal receptive field size that functions effectively across different scenarios is essential. Our experiments, detailed in Tab. 3(f)-(b), explore the effects of varying context filter sizes from 3 to 11. The optimal performance is achieved with a filter size of 9; larger sizes led to decreased performance, suggesting that excessively large receptive fields may detract from effective object matching by homogenizing the context features across all objects. Further investigation into different types of context filters shown in Tab. 3(f)-(d) reveals that the average filter, which evenly reflects surrounding information, offers a well-defined benefit. In contrast, the learnable filter, which lacks specific directives on characterizing surrounding information, performs similarly to scenarios without enhanced context features, as demonstrated in Tab. 3(f)-(a)-(v). This indicates the importance of a clearly defined context filter in improving the segmentation and tracking accuracy.
The number of adjacent frames used during training. Our fundamental assumption is that the surrounding information between adjacent frames remains relatively stable, thereby aiding object matching. Tab. 3(f)-(c) details the performance comparison based on the number of adjacent frames used during training. Utilizing three frames results in the highest performance, achieving 30.0 AP. Increasing the number of frames decreases performance, likely due to larger gaps between sampled frames which lead to more significant changes in the surrounding information, thus complicating object matching. Consequently, we have found that using three frames optimizes training effectiveness.
Context feature alignment. The tracking network aligns instance features and produces outputs in varying orders, necessitating the precise alignment of context features for accurate matching in subsequent frames. Misalignment of these features can lead to incorrect matching of context information for each object, significantly impacting performance. As demonstrated in Tab. 3(f)-(f), misalignment results in a notable performance drop of 3.3 AP. This underscores the critical need for accurate alignment of context information to ensure robust object tracking performance.
6 Conclusion
Acknowledgments
This work was supported by Korea Research Institute for defense Technology planning and advancement through Defense Innovation Vanguard Enterprise Project, funded by Defense Acquisition Program Administration(R230206), Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.2014-3-00123, Development of High Performance Visual BigData Discovery Platform for Large-Scale Realtime Data Analysis) and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00210908).
Appendix A Appendix
A.1 Limitation
A.2 Experimental Details
A.2.1 Datasets
Youtube-VIS 2019 and 2021 YouTube-VIS was introduced by Yang et al. in their pioneering study on the VIS task (Yang et al., 2019). This dataset comprises high-resolution YouTube videos, categorized into 40 distinct classes. The 2019 version of the dataset includes 2,238 videos for training, 302 for validation, and 343 for testing (Yang et al., 2019). The 2021 update expands these numbers to 2,985, 421, and 453 videos for training, validation, and testing, respectively (Yang et al., 2021a). YouTube-VIS is utilized across various pixel-level video understanding tasks, including VIS, video semantic segmentation, and video object detection.
OVIS The OVIS dataset (Qi et al., 2022) presents a significant challenge with its frequent occlusions and a realistic representation of common everyday objects. This makes it highly relevant for real-world applications. OVIS videos are longer and contain more objects compared to those in YouTube-VIS, which increases the complexity of segmentation and tracking tasks. The dataset is organized into training, validation, and test sets, with 607, 140, and 154 videos, respectively.
VIPSeg VIPSeg (Miao et al., 2022) is a comprehensive Video Panoptic Segmentation dataset that includes 3,536 videos and 84,750 frames, annotated with pixel-level panoptic labels. Unlike earlier VPS datasets that primarily focus on street views, VIPSeg offers a broader range of challenges and practical scenarios. It features 232 diverse settings and is annotated with 58 ‘thing’ classes and 66 ‘stuff’ classes, making it one of the most diverse and challenging datasets available in the field.
A.2.2 Implementation
Our segmentation approach employs the Mask2Former architecture (Cheng et al., 2022), utilizing the officially recommended hyperparameters. For all experimental settings, we follow established practices by incorporating COCO joint training, as adopted in previous methodologies (Wu et al., 2022b; Heo et al., 2022; 2023; Ying et al., 2023; Zhang et al., 2023a). The tracking network consists of six transformer blocks. Within the tracking network’s transformer blocks, we innovate by replacing the standard cross-attention layer with the referring cross-attention layer, as introduced in (Zhang et al., 2023a). Additionally, we conduct experiments with the temporal refiner (Zhang et al., 2023a) over 160k iterations, specifically analyzing sequences of 15 consecutive frames to enhance tracking accuracy.
For efficient training, we adopt a staged approach where the segmentation network is trained first, followed by the tracking network with all other parameters frozen, promoting stability and efficiency in learning, as suggested by previous studies (Zhang et al., 2023a; Li et al., 2023a). Optimization is carried out using the AdamW optimizer (Loshchilov & Hutter, 2017), with a starting learning rate of 1e-4 and a weight decay of 5e-2. The training process spans 40k iterations for the segmentation network and 160k iterations for the tracking network, with learning rate reductions scheduled at 28k and 112k iterations, respectively. During training, we sample three frames for the segmentation network and five frames for the tracking network from each of eight batched videos. These frames undergo resizing to ensure the shorter side is between 320 and 640 pixels, while the longer side does not exceed 768 pixels. The loss function weights are set to $\lambda_{\text{cls}}=2.0$, $\lambda_{\text{bce}}=5.0$, $\lambda_{\text{dice}}=5.0$, $\lambda_{\text{ctx}}=2.0$, and $\lambda_{\text{pro}}=2.0$ to balance the contributions of each component during training. For inference, the shorter side of input frames is scaled down to 448 pixels to maintain a consistent aspect ratio across inputs. All experiments are conducted using 8 RTX2080Ti GPUs for the ResNet-50 backbone and 8 RTX3090 GPUs for the Swin-L and ViT-L backbones, ensuring adequate computational resources are available for the demands of each model configuration.
A.3 Additional Experiments
Analysis on object embeddings. To demonstrate the effectiveness of our context-aware instance learning, we compare the distribution of object embeddings from three different models, as shown in Fig. 4. MinVIS does not engage in video learning, resulting in less effective distinction between objects. Compared to MinVIS, CTVIS shows a clearer object distinction by employing contrastive learning among object embeddings, but it still exhibits some overlaps in object clusters. In contrast, CAVIS forms much more distinct object clusters, highlighting the advantage of leveraging contextual information for object identification.
Figure 4: Visualization of object embeddings. Each point on the t-SNE (Van der Maaten & Hinton, 2008) plot represents the learned object embeddings. The three different colors of points indicate the embeddings of three different elephants throughout the entire video.
Qualitative results. We provide additional qualitative results of CAVIS across various datasets, as depicted in Fig. 5-8. These results underscore the robust capability of CAVIS to track objects in diverse scenarios for both VIS and VPS tasks. Notably, CAVIS excels in environments featuring numerous similar objects, fast-moving objects, and significant occlusions, demonstrating its effectiveness across complex dynamic scenes.

### v2

@@ 4.2 Prototypical Cross-frame Contrastive Loss
Current VIS methodologies emphasize the importance of object feature representation learning, especially in tracking tasks where matching object features across frames is crucial. In a query-based segmenter, object features $\hat{Q}_{t}^{n}\in\mathbb{R}^{C}$ are semantically correlated with each pixel embedding of the feature map $F_{t}^{\{h,w\}}\in\mathbb{R}^{C}$ from the pixel decoder, as they are used for mask prediction. This ensures consistent feature representation within object-containing regions and introduces an intra-frame constraint that reflects similarity within the feature map. By examining the inter-frame relationships of pixel embeddings, our method improves object association, essential for contrastive learning methods that strive to differentiate between similar and dissimilar objects across frames.
@@
$\begin{gathered}\mathcal{L}_{\text{PCC}}=\sum_{t=1}^{T}\sum_{n=1}^{N_{GT}}\mathcal{L}_{\text{Emb}}\left(\eta_{t}^{\hat{\sigma}^{t}(n)}\right),\\ \eta_{t}^{n}=\frac{\sum_{h=1}^{H}\sum_{w=1}^{W}F_{t}^{\{h,w\}}*\mathbbm{1}\left(M_{t}^{\{h,w\}}==1\right)}{\sum_{h=1}^{H}\sum_{w=1}^{W}\mathbbm{1}\left(M_{t}^{\{h,w\}}==1\right)}.\end{gathered}$ (13)
@@ 4.3 Training Loss
To train the tracking network $\mathcal{T}$, we calculate the matching cost only for objects that appear for the first time to ensure consistent video-level matching pairs, as implemented in prior work [57]. The objective function $\mathcal{L}_{\mathcal{T}}$ incorporating the matching relation $\hat{\sigma}_{\text{con}}$ is defined as follows:
$\begin{gathered}\mathcal{L}_{\mathcal{T}}=\sum_{t=1}^{T}\sum_{n=1}^{N_{GT}}\mathcal{L}_{\text{Inst}}\left(y^{t}_{n},\hat{y}^{t}_{\hat{\sigma}_{\text{con}}(n)}\right),\\ \hat{\sigma}_{\text{con}}=\underset{\sigma\in\mathfrak{S}_{N}}{\arg\min}\sum_{k=1}^{N_{GT}}\mathcal{L}_{\text{Match }}\left(y_{k}^{f(k)},\hat{y}_{\sigma(k)}^{f(k)}\right),\end{gathered}$ (15)
@@ 5 Experiments
We evaluate CAVIS on two major tasks: video instance segmentation (VIS) and video panoptic segmentation (VPS) on four benchmark datasets recognized for their challenges and prevalence in the research community: YouTubeVIS-2019 [51], YouTubeVIS-21 [52], OVIS [38], and VIPSeg [32]. For VIS, performance metrics include average precision (AP) and average recall (AR) as established in previous studies [51]. In the realm of VPS [21], we further examine our model’s capabilities using the Segmentation and Tracking Quality (STQ) metric, and the Video Panoptic Quality (VPQ) metric.
@@ 5.1 Implementation Details
We employ Mask2Former [10] as our segmentation network, utilizing three distinct backbone encoders: ResNet-50 [15], Swin-L [30], and ViT-L [13]. The ResNet-50 and Swin-L backbones are initialized with parameters pre-trained on the COCO dataset [29], while the ViT-L backbone uses initialization parameters from DINOv2 [37]. Additionally, for the ViT-L backbone, we employ a memory-efficient version of VIT-Adapter [7], aligning with recent advancements in network efficiency [58]. Further details are described in the Supplementary Material.
Table 3: Comparison on VIPSeg validation sets. ‘Th’ and ‘St’ denote ‘things’ and ‘stuff’ classes. denotes the model trained with the temporal refiner [57].
Method ResNet-50 ViT-L VPQ $\text{VPQ}^{\text{Th}}$ $\text{VPQ}^{\text{St}}$ STQ VPQ $\text{VPQ}^{\text{Th}}$ $\text{VPQ}^{\text{St}}$ STQ VPSNet-SiamTrack[45] 17.2 17.3 17.3 21.1 - - - - VIP-Deeplab[39] 16.0 12.3 18.2 22.0 - - - - Clip-PanoFCN[33] 22.9 25.0 20.8 31.5 - - - - Video K-Net [27] 26.1 - - 31.5 - - - - TarVIS[1] 33.5 39.2 28.5 43.1 - - - - Tube-Link[28] 39.2 - - 39.5 - - - - Video-kMax [41] 38.2 - - 39.9 - - - - DVIS [57] 39.4 38.6 40.1 36.3 - - - - DVIS++ [58] 41.9 41.0 42.7 38.5 56.0 58.0 54.3 49.8 Ours 42.4 43.1 41.8 39.7 56.9 60.1 54.2 51.0 DVIS [57] 43.2 43.6 42.8 42.8 - - - - DVIS++ [58] 44.2 44.5 43.9 43.6 58.0 61.2 55.2 56.0 Ours 45.3 47.5 43.4 45.3 58.5 63.1 54.5 56.1
@@ 5.3 Ablation Study
We conduct ablation studies on the OVIS dataset [38] with the ResNet-50 [15] backbone, detailed in Tab. 3(e). We use the same experimental setting as those in the main experiments. To evaluate the segmentation network across these setups in Tab. 3(e)-(a-c), we employ the minimal post-processing method proposed by MinVIS [19].
@@
Ablation study on technical contributions of CAVIS. We conduct a series of experiments in Tab. 3(e)-(a) to demonstrate the effectiveness of our key components: the context-aware instance tracker (CAIT) and prototypical cross-frame contrastive (PCC) loss.
@@
Tab. 3(e)-(a) includes six experiments (i- vi), where experiment (i) present our baseline performance of retraining MinVIS [19] and DVIS [57] to match our settings. For segmentation netowrk ($\mathcal{S}$), experiments (iii- iv) show that implementing contrastive learning with context-aware instance features results in a notable +3.3 AP improvement over the baseline MinVIS. Further performance boosts are noted when PCC is introduced in experiments (v- vi), which records the highest performance of 30.0 AP by utilizing both $\mathcal{L}_{\text{CTX}}$ and $\mathcal{L}_{\text{PCC}}$. This highlights the synergistic effect of integrating context-aware tracking with cross-frame contrastive loss, significantly enhancing the system’s accuracy and effectiveness.
@@
We also evaluate the tracking network ($\mathcal{T}$) by initially using each fixed pre-trained segmentation network listed in Tab. 3(e)-(i- vi). Comparing setups (v) and (vi), our newly designed context-aware cross-attention improves performance by +2.3 AP over the standard cross-attention, achieving 37.6 AP. These findings validate the efficacy of the context-aware feature, confirming its significant advantages for instance matching in complex video scenarios.
Table 4: Ablation studies on each component of CAVIS.
Segmenter ($\mathcal{S}$) Tracker ($\mathcal{T}$)
$\mathcal{L}_{\text{CTX}}$ $\mathcal{L}_{\text{PCC}}$ AP Context-aware matching AP
(i) 26.4 33.2
(ii) 28.1 34.2
(iii) 29.7 34.8
(iv) 29.7 37.2
(v) 30.0 35.3
(vi) 30.0 37.6
Metric: AP The numbe of frames
2 3 4
Filter size $3\times 3$ 27.3 27.3 27.2
$5\times 5$ 28.3 28.7 28.1
$7\times 7$ 28.7 29.3 28.1
$9\times 9$ 29.5 30.0 28.7
$11\times 1$ 11 28.7 29.1 28.0
Metric Context filter type in $\mathcal{S}$
Average Max Median Learnable
AP 30.0 29.4 29.6 28.7
Context alignment in $\mathcal{T}$
33.2 37.6
Value for CAIT
${Q}$ $\hat{Q}$
36.8 37.6
Context filter. Given that images feature objects at various scales, identifying the optimal receptive field size that functions effectively across different scenarios is essential. Our experiments, detailed in Tab. 3(e)-(b), explore the effects of varying context filter sizes from 3 to 11. The optimal performance is achieved with a filter size of 9; larger sizes led to decreased performance, suggesting that excessively large receptive fields may detract from effective object matching by homogenizing the context features across all objects. Extended analysis on this aspect is provided in our supplementary materials. Further investigation into different types of context filters shown in Tab. 3(e)-(c) reveals that the average filter, which evenly reflects surrounding information, offers a well-defined benefit. In contrast, the learnable filter, which lacks specific directives on characterizing surrounding information, performs similarly to scenarios without enhanced context features, as demonstrated in Tab. 3(e)-(a)-(ii). This indicates the importance of a clearly defined context filter in improving the segmentation and tracking accuracy.
The number of adjacent frames used during training. Our fundamental assumption is that the surrounding information between adjacent frames remains relatively stable, thereby aiding object matching. Tab. 3(e)-(b) details the performance comparison based on the number of adjacent frames used during training. Utilizing three frames results in the highest performance, achieving 30.0 AP. Increasing the number of frames decreases performance, likely due to larger gaps between sampled frames which lead to more significant changes in the surrounding information, thus complicating object matching. Consequently, we have found that using three frames optimizes training effectiveness.
Context feature alignment. The tracking network aligns instance features and produces outputs in varying orders, necessitating the precise alignment of context features for accurate matching in subsequent frames. Misalignment of these features can lead to incorrect matching of context information for each object, significantly impacting performance. As demonstrated in Tab. 3(e)-(d), misalignment results in a notable performance drop of 4.4 AP. This underscores the critical need for accurate alignment of context information to ensure robust object tracking performance.
Value for context-aware instance matching. Context-aware features ($Q$), which include information on instance features, could be used as the value in context-aware matching. However, during segmenter training, the instance features ($\hat{Q}$) specifically drive the segmentation prediction. Therefore, it is more effective to use instance features as the value for matching, leading to better performance, as shown in Tab. 3(e)-(e).
6 Conclusion
Acknowledgments
This work was supported by Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No.RS-2014-II140123, Development of High Performance Visual BigData Discovery Platform for Large-Scale Realtime Data Analysis), the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. RS-2023-00210908) and the Institute of Information & Communications Technology Planning & Evaluation(IITP) grant funded by the Korea government(MSIT) (No. RS-2025-02219277, AI Star Fellowship Support Project(DGIST)).
Appendix A Appendix
{strip}[Uncaptioned image]
Figure 4: Potential error due to inaccurate mask predictions from the segmentation network.
A.1 Limitation
Potential error in prediction. Our model is designed to improve tracking accuracy by achieving precise object matching across frames rather than focusing on segmentation performance. Consequently, if the pretrained segmentation network produces inaccurate segmentation results, performance may decrease. However, even in scenarios with imprecise mask predictions, our proposed context-aware modeling can robustly track objects, as demonstrated in Fig. 4.
Figure 5: Visualization of object embeddings. Each point on the t-SNE [43] plot represents the learned object embeddings. The three different colors of points indicate the embeddings of three different elephants throughout the entire video.
Figure 6: Comparison of VIS results for the video in Fig. 5. These results show that our model robustly tracks objects even in scenes with severe occlusions.
A.2 Experimental Details
A.2.1 Datasets
Youtube-VIS 2019 and 2021 YouTube-VIS was introduced by Yang et al. in their pioneering study on the VIS task [51]. This dataset comprises high-resolution YouTube videos, categorized into 40 distinct classes. The 2019 version of the dataset includes 2,238 videos for training, 302 for validation, and 343 for testing [51]. The 2021 update expands these numbers to 2,985, 421, and 453 videos for training, validation, and testing, respectively [52]. YouTube-VIS is utilized across various pixel-level video understanding tasks, including VIS, video semantic segmentation, and video object detection.
OVIS The OVIS dataset [38] presents a significant challenge with its frequent occlusions and a realistic representation of common everyday objects. This makes it highly relevant for real-world applications. OVIS videos are longer and contain more objects compared to those in YouTube-VIS, which increases the complexity of segmentation and tracking tasks. The dataset is organized into training, validation, and test sets, with 607, 140, and 154 videos, respectively.
VIPSeg VIPSeg [33] is a comprehensive Video Panoptic Segmentation dataset that includes 3,536 videos and 84,750 frames, annotated with pixel-level panoptic labels. Unlike earlier VPS datasets that primarily focus on street views, VIPSeg offers a broader range of challenges and practical scenarios. It features 232 diverse settings and is annotated with 58 ‘thing’ classes and 66 ‘stuff’ classes, making it one of the most diverse and challenging datasets available in the field.
Figure 7: VIS results with various filter sizes.
A.2.2 Implementation
Our segmentation approach employs the Mask2Former architecture [10], utilizing the officially recommended hyperparameters. For all experimental settings, we follow established practices by incorporating COCO joint training, as adopted in previous methodologies [46, 17, 18, 55, 57]. The tracking network consists of six transformer blocks. Within the tracking network’s transformer blocks, we innovate by replacing the standard cross-attention layer with the referring cross-attention layer, as introduced in [57]. Additionally, we conduct experiments with the temporal refiner [57] over 160k iterations, specifically analyzing sequences of 15 consecutive frames to enhance tracking accuracy.
For efficient training, we adopt a staged approach where the segmentation network is trained first, followed by the tracking network with all other parameters frozen, promoting stability and efficiency in learning, as suggested by previous studies [57, 25]. Optimization is carried out using the AdamW optimizer [31], with a starting learning rate of 1e-4 and a weight decay of 5e-2. The training process spans 40k iterations for the segmentation network and 160k iterations for the tracking network, with learning rate reductions scheduled at 28k and 112k iterations, respectively. During training, we sample three frames for the segmentation network and five frames for the tracking network from each of eight batched videos. These frames undergo resizing to ensure the shorter side is between 320 and 640 pixels, while the longer side does not exceed 768 pixels. The loss function weights are set to $\lambda_{\text{cls}}=2.0$, $\lambda_{\text{bce}}=5.0$, $\lambda_{\text{dice}}=5.0$, $\lambda_{\text{ctx}}=2.0$, and $\lambda_{\text{pro}}=2.0$ to balance the contributions of each component during training. For inference, the shorter side of input frames is scaled down to 448 pixels to maintain a consistent aspect ratio across inputs. All experiments are conducted using 8 RTX2080Ti GPUs for the ResNet-50 backbone and 8 RTX3090 GPUs for the Swin-L and ViT-L backbones, ensuring adequate computational resources are available for the demands of each model configuration.
A.3 Further Studies
Analysis on object embeddings. To demonstrate the effectiveness of our context-aware instance learning, we compare the distribution of object embeddings from three different models, as shown in Fig. 5. MinVIS does not engage in video learning, resulting in less effective distinction between objects. Compared to MinVIS, CTVIS shows a clearer object distinction by employing contrastive learning among object embeddings, but it still exhibits some overlaps in object clusters. In contrast, CAVIS forms much more distinct object clusters, highlighting the advantage of leveraging contextual information for object identification. This trends are reflected in the VIS results, as shown in Fig. 6.
@@ A.3 Further Studies
Metric Context alignment
@@
Figure 8: VIS results from our model on a video containing a fast-moving object.
Effective filter size. Videos often contain objects of varying sizes, and for smaller objects, using an excessively large context area can introduce noise, leading to inaccurate matching as shown in Fig. 7. To better understand this effect, we analyze the impact of different filter sizes to identify the optimal value. Our findings indicate that the overall trend remains consistent, regardless of variations in the number of frames used during training.
Ablation study with minimal setups. To simplify reproducibility, we additionally provide ablation studies on the OVIS dataset [38] with the ResNet-50 [15] backbone, detailed in Tab. 7(f). The results exhibit similar trends to those observed in the main text, further validating the consistency of our findings. For these experiments, we train the segmentation network with 2 frames over 40k iterations, while the tracking network is trained with 5 frames over 40k iterations. Experiments (i- iii) show that implementing contrastive learning, whether with standard or context-aware instance features, leads to significant performance gains. Particularly, context-aware instance features result in a notable +2.7 AP improvement over the baseline, a considerable increase compared to the +1.3 AP improvement observed with standard instance features.
Robustness of our model. Our method does not rely solely on context. By incorporating both context and instance features, our approach shows robustness even in scenes containing fast-moving objects where context changes rapidly, as shown in Fig. 8.
Method AP
MinVIS [19] 23.3
DVIS [57] 31.6
VITA [17] 32.6
DVIS++ [58] 37.2
GenVIS [18] 37.5
Ours 38.6
Table 6: Comparison on YTVIS 2022 dataset.
Performance on long video. We additionally report the performance on the YouTube-VIS 2022 dataset, a well-known benchmark featuring long video sequences. Its validation set includes 71 additional videos compared to the YouTube-VIS 2021 dataset, making it particularly challenging due to the need for accurately tracking dynamically appearing and disappearing objects over extended periods. We evaluate our model on these 71 long videos and compare it against existing state-of-the-art models with a ResNet-50 backbone. As shown in Tab. 6, our approach outperforms existing methods, demonstrating that our context-aware modeling remains effective for robust object matching even in long-range video scenarios.
Method Time (ms) YTVIS19 OVIS
DVIS [57] 78.9 51.2 30.2
GenVIS [18] 80.1 50.0 35.8
Ours 85.6 55.7 37.6
Table 7: Computational cost.
@@
Computational cost. We compare the inference speed of our approach against recent state-of-the-art methods, DVIS [57] and GenVIS [18], as shown in Tab. 7. The inference speeds were measured under identical conditions on a 2080ti GPU using the ResNet-50 backbone. Our method requires an additional time cost of 6.7ms and 5.5ms compared to DVIS and GenVIS, respectively. However, this cost is justified by performance gains of +1.8 AP and +7.6 AP on OVIS, and +4.5 AP and +5.7 AP on YTVIS19, demonstrating a reasonable trade-off between increased computation and improved accuracy.
Additional qualitative results. We provide additional qualitative results of CAVIS across various datasets, as depicted in Fig. 9-12. These results underscore the robust capability of CAVIS to track objects in diverse scenarios for both VIS and VPS tasks. Notably, CAVIS excels in environments featuring numerous similar objects, fast-moving objects, and significant occlusions, demonstrating its effectiveness across complex dynamic scenes.